How much 0.1 F HC2H3O2 is required to make 500ml of solution with a pH of 4.0?


1
Expert's answer
2022-03-13T15:20:47-0400

HC2H3O2 is the acetic acid - CH3COOH;

[H+] = 10-pH = 10-4 = 0,0001М;

pKa(CH3COOH) = 4.75;

Ka(CH3COOH) = 10-pKa(CH3COOH) = 10-4.75 = 1.8 * 10-5;

CH3COOH = H+ + CH3COO-

x 0 0

-0.0001 0.0001 0.0001

x-0.0001 0.0001 0.0001

Ka = [H+] * [CH3COO-]/[CH3COOH];

Ka = 0.0001 * 0.0001/(x-0.0001);

1.8*10-5 = 1*10-8/(x-0.0001);

x-0.0001 = 1*10-8/1.8*10-5 = 5.56 * 10-2;

x = 5.56 * 10-2 + 0.0001 = 5.57 * 10-2;

C1 * V1 = C2 * V2;

5.57 * 10-2 * 500 = 0.1 * Y;

Y = 5.57 * 10-2 * 500/0.1 = 278.5 mL.

Answer: 278.5 mL.